you have been given the mathematical model to calculate velocity of a car accelerating from rest in straight line: v(t) = A (1-e ^ - t/tmaxspeed) (anything after the symbol ^ is to the power of) v(t is instantaneous velocity of car (m/s), t is time in seconds, tmaxspeed is time to reach maximum speed in seconds and A is constant. you are given the information that t (0-28 m/s) is 2.6s, t (400m) is 10.46s & tmaxspeed is 7s. please derive an equation a(t) for the instantaneous position of the car as a function of time? Identify the acceleration of the car at t = 0s and asymptote of this function as t → ∞?
1
Expert's answer
2020-02-27T09:21:08-0500

"v(t)=A(1\u2212e^{\u2212t\/t_{maxspeed}})"

using the information that t (0-28 m/s) is 2.6s & tmaxspeed is 7s

"v(2.6)=28=A(1-e^{-2.6\/7}) => A=90.25m\/s"

finally "v(t)=90.25(1-e^{-t\/7})"


Acceleration: "acc(t)=dv\/dt=90.25\/7*e^{-t\/7}=12.9e^{-t\/7}"

"acc(t=0)=12.9m\/s^2"

Asymptote of this function as t → ∞:

"lim_{t \to \infty}acc(t)=lim_{t \to \infty}12.9e^{-t/7}=0" ,

then we have horizontal asymptote as t → ∞ : "y(t)=0."


Derive an equation a(t) for the instantaneous position of the car as a function of time:

"a(t)=a_0+A*t+A*t_{maxspeed}*e^{-t\/t_{maxspeed}}"

using that t(400m) is 10.46s & tmaxspeed is 7s & A=90.25:

"a(10.46)=400=a_0+90.25*10.46+90.25*7*e^{-10.46\/7}"

"=> a_0=-685.8m"

finally: "a(t)=-685.8+90.25*t+90.25*7*e^{-t\/7}"